If only digits 0,1,2,3,4,5,6 maybe used, find the number of possibilities in each category:
a) three-digit numbers
b) odd three-digit numbers
c) three-digit numbers without repeated digitits

Answer by ankor@dixie-net.com(22740)

If only digits 0,1,2,3,4,5,6 maybe used, find the number of possibilities in each category:
:
Assuming the 1st digit cannot be 0, first digit has 6 choices
a) three-digit numbers, digits can be repeated
6*7*7 = 294
:
b) odd three-digit numbers
6*7*3 = 126
:
c) three-digit numbers without repeated digitits
6*6*5 = 180
:
Just what is "digitits" something to do with breast implants, perhaps?